The singles court in badminton is 27 feet longer than it is wide.If its perimeter is 122 feet, what are the length and the width.
Answer by macston(5194)

.
W=width; L=length=W+27ft
.
P=2(L+W)
122ft=2(L+W)
61ft=L+W
61ft=(W+27ft)+W
61ft=2W+27ft
34ft=2W
17ft=W
.
ANSWER 1: The width is 17 feet.
.
L=W+27ft=17ft+27ft=44ft
.
ANSWER 2: The length is 44 feet.
.
CHECK:
P=2(L+W)
122ft=2(44ft+17ft)
122ft=2(61ft)
122ft=122ft
.